OpenStudy (anonymous):

Graph using the slope and y-intercept 2x+5y=-25

OpenStudy (wolfe8):

Can you distinguish and obtain slope and y-intercept from functions? Hint: y=mx+b

OpenStudy (anonymous):

I have solved it before am just having problem in the graph ,are you good in graph

OpenStudy (wolfe8):

I wouldn't say I'm good but I think I can manage this one.

OpenStudy (jdoe0001):

so what would you get if you solve for "y"?

OpenStudy (anonymous):

that was my final answer Y=-5(2/5)x

OpenStudy (jdoe0001):

2x+5y=-25 # subtracting -2x from each side 5y = -25 - 2x # dividing by 5 each side \(\bf 5y = -25 - 2x \implies y = -\cfrac{25}{5}-\cfrac{2x}{5}\implies y =-\cfrac{2}{5}x-5 \)

OpenStudy (anonymous):

PLOT THE GRAPH